OpenStudy (anonymous):

add ..-(5x-9y)/(8x)x(9x-6y)/(8x)

OpenStudy (anonymous):

\[\frac{-(5x+9)}{8x}+\frac{9x-6y}{8x}\] or \[\frac{-(5x-9y)}{8x}*\frac{9x-6y}{8x}?\]

OpenStudy (radar):

\[4x-6y-9\over8x\]